(a) The graph of y=f(x) is shown. Translate it to get the graph of y=f(x+5).
(b) The graph of y=g(x) is shown. Translate it to get the graph of y=g(x)-2.

Answer by ikleyn(52781)

.
(a) The graph of y=f(x) is shown. Translate it to get the graph of y=f(x+5).
(b) The graph of y=g(x) is shown. Translate it to get the graph of y=g(x)-2.
~~~~~~~~~~~~~~~~

(a)  To get the graph y = f(x+5), translate the graph y = f(x) horizontally 
     5 units to the left.


(a)  To get the graph y = g(x)-2, translate the graph y = g(x) vertically 
     2 units down.

Solved.